# Bipartite Graph Variational Auto-Encoder with Fair Latent Representation to Account for Sampling Bias in Ecological Networks

### v2

@@ Abstract
We propose a method to represent bipartite ecological networks tailored to tackle sampling bias associated with accumulating biodiversity data coming from citizen science based monitoring schemes. We adapt the variational graph auto-encoder to the bipartite case to generate embeddings where the two sets of nodes are positioned based on their probability of connection. We translate the fairness framework commonly considered in sociology in order to address sampling bias in this ecological dataset. By incorporating the Hilbert-Schmidt independence criterion, we ensure that the embedding is independent of continuous variables related to the sampling process.
Keywords: Citizen Science, Ecological network, Graph Neural Network, Hilbert-Schmidt Independence Criterion, Sampling Effect
@@ 1 Introduction
Citizen science programs facilitate the accumulation of biological or ecological data and are progressively more prevalent in biodiversity monitoring efforts (Conrad and Hilchey, 2011; Chandler et al., 2017; Pocock et al., 2018). However, these methods of collecting data are prone to sampling bias due to the multiplication of observers and associated observer effects (Bird et al., 2014). Further, one notable bias encountered in citizen science programs is the progressive accumulation of user experience in data collection (Kelling et al., 2015). For example, Jiguet (2009) found in a breeding bird survey that observers tended to count more birds after the first year compared to the initial year of observation. In order to mitigate the bias introduced by the observers, Johnston et al. (2018) suggested estimating the observers’ expertise and incorporate it as a covariate in their model.
As a study case, we focus on a citizen science based monitoring schemes where participants collect data on plant pollinator interactions across France. These data allow an unprecedented understanding of the variations of plant-pollinator ecological networks (Deguines et al., 2012; Pocock et al., 2016). As other dataset collected using citizen science, the Spipoll dataset is prone to bias caused by variations in users skills and experience. Given the amount of observed interactions and the need for bias mitigation, network analysis on this very large plant-pollinator network could profit from recent developments in machine learning such as graph embeddings using graph neural networks (GNN) and a dedicated bias mitigation strategy. Here, we propose a framework to reconstruct plant-pollinator network from such citizen science biodiversity monitoring schemes while accounting for associated sampling bias.
@@
One noticeable bias in the Spipoll dataset is that observers tend to record more pollinating insects per observation session as they get more experienced. This is illustrated in Figure 1 where the average number of observed insects per session increases with the number of time the user has performed the protocol, with different slopes depending on the insect order. Such bias is related to both the ability of observers to differentiate and identify the various insect species which has been shown to increase with observer experience Deguines et al. (2018) and the increasing photographing skills that allow the observers not to miss small and/or fast moving pollinators. In this context, Deguines et al. (2016, 2018) addressed the potential variation in the number of photographed insects among different observers by incorporating observer identity as random terms in their models.
@@
Graph embedding regroups different methods, allowing to represent a network into a vector space in order to gain understanding of key network features. These methods are especially important in the context of large networks. Recently developed graph neural networks (GNNs) enable graph embedding with large-scale methods such as graph isomorphism network (Xu et al., 2019), graph attention network (Veličković et al., 2018) or the variational graph auto-encoder (Kipf and Welling, 2016). All these methods can also handle numerous covariates on nodes.
GNNs are currently growing in popularity in various domains such as bioinformatics (Zhang et al., 2021) and chemistry (Reiser et al., 2022). In ecology, networks have been analyzed to study various types of interaction between species of plants and animals (Ings, 2009). The stochastic block model (Nowicki and Snijders, 2001) and the latent block model (Govaert and Nadif, 2010) for bipartite graphs are notorious models using latent variable in ecology (Terry and Lewis, 2020; Durand-Bessart et al., 2023). While graph embedding methods start being used for ecological networks (e.g. Botella et al. (2022), Strydom et al. (2022)), GNNs have yet to be diffused in that research field. GNNs could be particularly relevant for ecological networks because very large data sets of interactions among species are now becoming available (e.g. through the development of citizen science programs), in addition to many covariates linked to the nodes (e.g. species name and traits, environmental characteristics at the time of interaction observation). An important issue with the analysis of ecological networks is related with the strong effects of sampling effort and methods on network structure (Jordano, 2016; Doré et al., 2021). Sampling interactions among species to build ecological networks is indeed a daunting task. For example, extensive sampling of plant-pollinator interactions has been shown to only provide a subset of the existing interactions (Chacoff et al., 2012; Jordano, 2016) and sampling protocols are known to be subject to bias (Gibson et al., 2011), both of which need to be accounted for to make meaningful ecological interpretations (Doré et al., 2021). One could wish to have an embedding which is independent of a certain set of covariates linked to such sampling effects and related bias. This can be of particular interest for citizen science programs, where biases can arise from the observer’s experience level (Jiguet, 2009; Deguines et al., 2018).
@@
The solution to correct for sampling biases in ecological network analysis may be found in machine learning applied to sociology with the notion of fairness (Caton and Haas, 2020; Carey and Wu, 2022), where the main idea is to train an algorithm whose predictions are independent of a protected variable. It is used in sociology (Carey and Wu, 2022) to have predictions that are independent of gender, sexuality or disability. Recently, works on fairness have been extended to social network analysis (Saxena et al., 2022), for fair link prediction (Li et al., 2021), fair graph exploration (Rahman et al., 2019). Fairness has also been developed for variational auto-encoder, with the fair auto-encoder (Louizos et al., 2017) or with adversarial debiasing (Zhang et al., 2018), which can help to aim for a fair latent representation of the network. Most papers about fairness for networks seek for fairness regarding a binary (or a categorical) protected variable (Saxena et al., 2022), which is not always appropriate to the study of ecological networks where sampling biases might also be measured by continuous variables (e.g. sampling time, observer’s experience level). The adversarial debiasing from Zhang et al. (2018) could be applied to continuous variables, but has theoretical guarantees for the discrete protected variable. Even if the fair variational auto-encoder (Louizos et al., 2017) assumes independent prior, they penalize their loss with a regularization term that considers a discrete protected variable. In recommender systems (Beel et al., 2016; Li et al., 2023) where bipartite networks are also encountered, different notions of fairness for different characteristics on graphs with heterogeneous nodes (Li et al., 2023) are studied but the protected variable is binary or categorical. However, the targeted fairness which is relevant in recommender system, is not adapted to address sample bias in ecology since the former focuses on ensuring equitable recommendations for users (Li et al., 2023), which is a user-centric goal, while the latter deals with a broader scope of the whole network.
@@
There are other machine learning methods related with debiasing taking into account continuous variables, such as disentanglement methods. For example the $\beta-VAE$ (Higgins et al., 2016) can provide a latent space with some conditionally independent latent factors. However their goal is different because they do not seek complete independence between the latent representation and the sensitive factors (Locatello et al., 2019). The method we propose relies on the Hilbert-Schmidt independence criterion (HSIC) Gretton et al. (2005). The HSIC is a measure of independence between two random variables and can be used as a penalty term to ensure independence between continuous variable. This metric has been put in practice in learning context by Greenfeld and Shalit (2020), Ma et al. (2020), Pérez-Suay et al. (2017) or Quadrianto et al. (2019).
@@
In this paper, we aim to obtain a latent representation of the ecological network derived from Spipoll that is independent of covariates that may be continuous. First, we specify our goal in a simple context where linear embeddings are considered with a simple data structure. After presenting the necessary background on GNNs and the HSIC, we introduce the model that makes recourse to an extension of the variational graph auto-encoder to bipartite networks with a penalization based on the HSIC. We attest the performance of our proposed model with simulated data mimicking the context of Spipoll. Finally, we apply our methodology to the Spipoll data set , where we show that accounting for the sampling effort can change the understanding of the network.
2 Background
2.1 The linear embedding case
It is difficult to consider a fairness problem in the network framework because of the structure of the data itself. Ensuring fairness in network analysis involves addressing disparities in the representation and impact of the connection pattern and the covariates available. To have a better understanding of the problem at stake, we first present the available tools when we consider a linear embedding for a dataset living in a real linear space.
Let $X$ a $n\times d$ matrix and let $S$ be $n\times d_{s}$ matrix. Without loss of generality, we assume that each column of $X$ has been centered. We wish to perform a one dimensional principal component analysis on $X$ that would yield us a vector $v$ and a lower dimensional embedding of $X$ given by $Xv$ that maximizes the variance. However, we wish to have a latent representation $Xv$ independent of the protected variable $S$. If we were in the context of probabilistic PCA (Tipping and Bishop, 1999) where $X$ and $S$ would have been multivariate Gaussian, projecting $X$ onto the space orthogonal to $S$: $S^{\perp}$ beforehand would have been enough to guarantee the independence between $S$ and the latent representation $Xv$, this can be solved using PPCA with covariates (Kalaitzis and Lawrence, 2012). We show that this approach is equivalent to find the optimal projection with respect to an independence constraint.
Proposition 1
$L(v)=v^{\top}\Lambda v+\lambda_{1}(1-v^{\top}v)+\lambda_{2}||S^{\top}Xv||^{2}$ (1)
can also be obtained by computing the first component of the PCA of $P_{S^{\bot}}X$. The proof is provided in the supplementary material.
We have reformulated the linear embedding problem under independence constraint as an optimization problem with a supplementary term that encourages the independence. This optimization formulation can be used for graph network analysis. Embedding the graph as a vector using a variational graph auto-encoder (VGAE) (Kipf and Welling, 2016) would yield a Gaussian latent representation $Z$. As we cannot guarantee that $Z$ and $S$ are jointly Gaussian, we will use another criterion than the covariance to have independence between $Z$ and $S$: the Hilbert-Schmidt independence criterion (HSIC), first proposed by Gretton et al. (2005) which is a metric testing for the independence of two variables. Compared to the other proposed methods of embedding, the probabilistic setting of the GVAE fits well with the use of the HSIC, and its generative aspect allows network generation for various ecological contexts.
@@ 2.2 Bipartite variational graph auto-encoder
We adapt the variational graph auto-encoder from Kipf and Welling (2016) to the bipartite case by considering two graph convolutional networks (GCN), one for each node type. We consider a biadjacency matrix $B_{i,j}$ of size $n_{1}\times n_{2}$ representing our graph. Let
@@ 2.2.2 Decoder
Following Rubin-Delanchy et al. (2021), we decide to use as a decoder the generalised random dot product
@@
$p(B_{i,j}|Z_{1i},Z_{2j})=sigmoid(Z_{1i}^{\top}\mathbf{I}_{D_{+},D_{-}}Z_{2j})$ where $sigmoid:x\mapsto\frac{1}{1+e^{-x}}$ and $\mathbf{I}_{D_{+},D_{-}}$ is a diagonal matrix with $D_{+}$ ones followed by $D_{-}$ minus ones on its diagonal, such as $D_{+}+D_{-}=D$. The full auto-encoder can be summarized as
@@ 2.3.1 Definition
Let $X$ and $Y$ two random variables in $\mathcal{X}$ and $\mathcal{Y}$ and $(X,Y)$ be the joint probability distribution. Let $\mathcal{F}$ and $\mathcal{G}$ be the RKHS on $\mathcal{X}$ and $\mathcal{Y}$ with their associated kernel $K:\mathcal{X}\times\mathcal{X}\to\mathbb{R}$ and $L:\mathcal{Y}\times\mathcal{Y}\to\mathbb{R}$. Gretton et al. (2005) define the HSIC as the norm of the cross-variance operator between the distribution in the RKHS:
$HSIC(X,Y)$ $=||C_{X,Y}||^{2}$
@@
Using some specific kernels such as the Gaussian kernel $K(x_{i},x_{j})=e^{-\frac{||x_{i}-x_{j}||^{2}}{2\sigma^{2}}}$ it can be shown that $HSIC(X,Y)=0\iff X\bot Y$(Gretton et al., 2005).
@@ 2.3.2 Estimation
Under the assumption that $X$ and $Y$ are independent, it has been proved that the distribution of $n\times\widehat{HSIC}$ can be asymptotically approximated by a Gamma distribution (Gretton et al., 2007):
@@
During the learning phase, using the HSIC as a penalty between the coordinates in the latent space and the protected variable will assure that the latent space is as much as possible independent of the protected variable. At the end of the learning, we can check the independence by calculating the p-value of the test and compare it to the desired significance level. An example in the linear case is available in Section B.4.
@@ 2.4 Estimation in high dimension
Calculating $\widehat{HSIC}$ requires to compute the $n\times n$ Gram matrix and can be time-consuming for large value of $n$. As a substitute, we can use Random Fourier Features (RFF) (Rahimi and Recht, 2007). RFF for learning has been used in (Louizos et al., 2017) to minimize the maximum mean discrepancy (MMD) between two distributions, in a setting where the protected variable have two possible values. Using the Gaussian kernel, minimizing the MMD is equivalent to matching all the moments of the two distribution, making the outcome of the encoder fair. In our settings, the HSIC can be seen as the MMD between the joint distribution $(X,Y)$ and the product of the distribution $X$ and $Y$. If the joint distribution is close to the product of the two distribution then the $X$ and $Y$ will behave as they are independent.
Assume that $K(x_{i},x_{j})=e^{-\frac{1}{2}||x_{i}-x_{j}||^{2}}$ ($\sigma^{2}=1$), with $x_{i}\in\mathbb{R}^{d}$, let $D<n$, $\omega$ be a $D\times d$ matrix where all entries are independently drawn from $\mathcal{N}(0,1)$ and $b$ be $d$-dimensional vector with each entry independently drawn from $Unif([0,2\pi])$. For any $x_{i}$, its RFF is defined as
$\psi_{X}(x_{i})=\sqrt{\frac{2}{D}}cos(\omega x_{i}^{\top}+b)\in\mathbb{R}^{D}.$
The main idea behind this RFF representation is that $K(x_{i},x_{j})\approx\langle\psi_{X}(x_{i}),\psi_{X}(x_{j})\rangle$, this property will allow the computation to be much faster (see Section B.5), while having a small error term (Sutherland and Schneider, 2015). If we also define $\psi_{Y}$ the as RFF of $Y$, then finally we can estimate the RFF HSIC (Zhang et al., 2018) as done in Quadrianto et al. (2019):
$RFF\>HSIC=$ $\frac{1}{n^{2}}\left|\left|\sum_{i=1}^{n}\psi_{X}(x_{i})\psi_{Y}(y_{i})^{\top}\right.\right.\left.\left.-\frac{1}{n}\left(\sum_{i=1}^{n}\psi_{X}(x_{i})\right)\left(\sum_{i=1}^{n}\psi_{Y}(y_{i})\right)^{\top}\right|\right|^{2}_{2}.$ (3)
The RFF HSIC can then be computed accurately with complexity $O(D^{2})$. Several other methods of estimation for HSIC exists for large-scale problem (Zhang et al., 2018). For the rest of this paper, $D=100$ will provide enough accuracy.
3 Bipartite and fair graph variational auto-encoder
Suppose our goal is to construct the latent representation $Z_{1}$ of the bipartite variational auto-encoder such as it is independent of a protected variable denoted by $S$. We optimize our parameters $(W_{v,i})_{1\leq v,i\leq 2}$ as defined in Section 2.2.1 to minimize a compromise between the variationnal lower bound of the auto-encoder and the HSIC between $\mu_{1}$ and the protected variable $S$. Even if the reconstruction would be penalized, this would yield a latent-space $\tilde{Z_{1}}\sim\mathcal{N}(\mu_{1},diag(\sigma^{2}_{1}))$ independent of $S$. The complete loss of this auto-encoder can be written as:

### v3

@@ Abstract
Citizen science monitoring programs can generate large amounts of valuable data, but are often affected by sampling bias. We focus on a citizen science initiative that records plant-pollinator interactions, with the goal of learning embeddings that summarize the observed interactions while accounting for such bias. In our approach, plant and pollinator species are embedded based on their probability of interaction. These embeddings are derived using an adaptation of variational graph autoencoders for bipartite graphs. To mitigate the influence of sampling bias, we incorporate the Hilbert-Schmidt Independence Criterion (HSIC) to ensure independence from continuous variables related to the sampling process. This allows us to integrate a fairness perspective—commonly explored in the social sciences—into the analysis of ecological data. We validate our method through a simulation study replicating key aspects of the sampling process and demonstrate its applicability and effectiveness using the Spipoll dataset.
Keywords: Graph Neural Network, Hilbert-Schmidt Independence Criterion, Ecological network, Citizen Science, Sampling Effect
@@ 1 Introduction
Citizen science programs facilitate the accumulation of biological or ecological data and are progressively more prevalent in biodiversity monitoring efforts (conradReviewCitizenScience2011; chandlerContributionCitizenScience2017; pocockVisionGlobalBiodiversity2018a). However, these methods of collecting data are prone to sampling bias due to the multiplication of observers and associated observer effects (birdStatisticalSolutionsError2014). Further, one notable bias encountered in citizen science programs is the progressive accumulation of user experience in data collection (kellingCanObservationSkills2015). For example, jiguetMethodLearningCaused2009 found in a breeding bird survey that observers tended to count more birds after the first year compared to the initial year of observation. In order to mitigate the bias introduced by the observers, johnstonEstimatesObserverExpertise2018 suggested estimating the observers’ expertise and incorporate it as a covariate in their model.
As a study case, we focus on a citizen science based monitoring schemes, named Spipoll, where participants collect data on plant pollinator interactions across France. These data allow an unprecedented understanding of the variations of plant-pollinator ecological networks (deguinesWhereaboutsFlowerVisitors2012; pocockVisualisation2016).
As other dataset collected using citizen science, the Spipoll dataset is prone to bias caused by variations in users skills and experience.
Given the amount of observed interactions and the need for bias mitigation, network analysis on this very large plant-pollinator network could profit from recent developments in machine learning such as graph embeddings using graph neural networks (GNN) and a dedicated bias mitigation strategy.
Here, we propose a framework to reconstruct plant-pollinator network from such citizen science biodiversity monitoring schemes while accounting for associated sampling bias.
@@
One noticeable bias in the Spipoll dataset is that observers tend to record more pollinating insects per observation session as they get more experienced. This is illustrated in Figure 1 where the average number of observed insects per session increases with the number of time the user has performed the protocol, with different slopes depending on the insect order. Such bias is related to both the ability of observers to differentiate and identify the various insect species which has been shown to increase with observer experience deguinesFosteringCloseEncounters2018 and the increasing photographing skills that allow the observers not to miss small and/or fast moving pollinators. In this context, deguinesFunctionalHomogenizationFlower2016; deguinesFosteringCloseEncounters2018 addressed the potential variation in the number of photographed insects among different observers by incorporating observer identity as random terms in their models.
@@
Graph embedding regroups different methods, allowing to represent a network into a vector space in order to gain understanding of key network features. These methods are especially important in the context of large networks. Recently developed graph neural networks (GNNs) enable graph embedding with large-scale methods such as graph isomorphism network (xuHowPowerfulAre2019), graph attention network (velickovicGraphAttentionNetworks2018) or the variational graph auto-encoder (Kipf). All these methods can also handle numerous covariates on nodes.
GNNs are currently growing in popularity in various domains such as bioinformatics (zhangGraphNeuralNetworks2021) and chemistry (reiserGraphNeuralNetworks2022). In ecology, networks have been analyzed to study various types of interaction between species of plants and animals (ings_review_2009). The stochastic block model (nowickiEstimationPredictionStochastic2001) and the latent block model (govaert2010latent) for bipartite graphs are notorious models using latent variable in ecology (terry_finding_2020; durand-bessart_trait_2023). While graph embedding methods start being used for ecological networks (e.g. botellaAppraisalGraphEmbeddings2022, strydomFoodWebReconstruction2022), GNNs have yet to be diffused in that research field. GNNs could be particularly relevant for ecological networks because very large data sets of interactions among species are now becoming available (e.g. through the development of citizen science programs), in addition to many covariates linked to the nodes (e.g. species name and traits, environmental characteristics at the time of interaction observation). An important issue with the analysis of ecological networks is related with the strong effects of sampling effort and methods on network structure (jordano_sampling_2016; dore_relative_2021). Sampling interactions among species to build ecological networks is indeed a daunting task. For example, extensive sampling of plant-pollinator interactions has been shown to only provide a subset of the existing interactions (chacoff_evaluating_2012; jordano_sampling_2016) and sampling protocols are known to be subject to bias (GibsonSampling2011), both of which need to be accounted for to make meaningful ecological interpretations (dore_relative_2021). One could wish to have an embedding which is independent of a certain set of covariates linked to such sampling effects and related bias. This can be of particular interest for citizen science programs, where biases can arise from the observer’s experience level (jiguetMethodLearningCaused2009; deguinesFosteringCloseEncounters2018).
@@
The solution to correct for sampling biases in ecological network analysis may be found in machine learning applied to sociology with the notion of fairness (catonFairnessMachineLearning2020; careyFairnessFieldGuide2022), where the main idea is to train an algorithm whose predictions are independent of a protected variable. It is used in sociology (careyFairnessFieldGuide2022) to have predictions that are independent of gender, sexuality or disability. Recently, works on fairness have been extended to social network analysis (saxenaFairSNAAlgorithmicFairness2022), for fair link prediction (liDYADICFAIRNESSEXPLORING2021), fair graph exploration (rahmanFairwalkFairGraph2019). Fairness has also been developed for variational auto-encoder, with the fair auto-encoder (louizosVariationalFairAutoencoder2017) or with adversarial debiasing (zhangMitigatingUnwantedBiases2018), which can help to aim for a fair latent representation of the network. Most papers about fairness for networks seek for fairness regarding a binary (or a categorical) protected variable (saxenaFairSNAAlgorithmicFairness2022), which is not always appropriate to the study of ecological networks where sampling biases might also be measured by continuous variables (e.g. sampling time, observer’s experience level). The adversarial debiasing from zhangMitigatingUnwantedBiases2018 could be applied to continuous variables, but has theoretical guarantees for the discrete protected variable. Even if the fair variational auto-encoder (louizosVariationalFairAutoencoder2017) assumes independent prior, they penalize their loss with a regularization term that considers a discrete protected variable. In recommender systems (beelResearchpaperRecommenderSystems2016; li2023recent) where bipartite networks are also encountered, different notions of fairness for different characteristics on graphs with heterogeneous nodes (liFairnessRecommendationFoundations2023) are studied but the protected variable is binary or categorical. However, the targeted fairness which is relevant in recommender system, is not adapted to address sample bias in ecology since the former focuses on ensuring equitable recommendations for users (liFairnessRecommendationFoundations2023), which is a user-centric goal, while the latter deals with a broader scope of the whole network.
@@
There are other machine learning methods related with debiasing taking into account continuous variables, such as disentanglement methods. For example the $\beta-VAE$ (Higgins2016betaVAELB) can provide a latent space with some conditionally independent latent factors. However their goal is different because they do not seek complete independence between the latent representation and the sensitive factors (locatello2019fairness). The method we propose relies on the Hilbert-Schmidt independence criterion (HSIC, grettonMeasuringStatisticalDependence2005a). The HSIC is a measure of independence between two random variables and can be used as a penalty term to ensure independence between continuous variable. This metric has been put in practice in learning context by greenfeldRobustLearningHilbertSchmidt2020, maHSICBottleneckDeep2020, perez-suayFairKernelLearning2017 or quadriantoDiscoveringFairRepresentations2019.
@@
In this paper, we aim to obtain a latent representation of the ecological network derived from Spipoll that is independent of covariates that may be continuous. After presenting the necessary background on GNNs and the HSIC, we introduce the model that makes recourse to an extension of the variational graph auto-encoder to bipartite networks with a penalization based on the HSIC. We attest the performance of our proposed model with simulated data mimicking the context of Spipoll. Finally, we apply our methodology to the Spipoll data set , where we show that accounting for the sampling effort can change the understanding of the network.
2 Bipartite and fair variational auto-encoder
Our goal is to obtain an embedding of the nodes of the network. Since data collection is prone to sampling bias, our aim is to constrain the embedding to be independent of some variables related to the sampling condition. These variables correspond to what are called protected variables in a fairness paradigm. In the following, we present an extension of the variational graph auto-encoder (VGAE) (Kipf) to a bipartite network. Then, since ensuring fairness in the network context is difficult because it involves addressing disparities in the representation and impact of the connection pattern, we rely on the optimization of a variational loss that incorporates a statistical measure of dependence between random variables, namely the HSIC (grettonMeasuringStatisticalDependence2005a). The HSIC and its efficient computation in high dimension are exposed in the following subsection. Finally, the optimized loss to ensure the embedding with the indepence constrain is presented in the final subsection.
@@ 2.1 Bipartite variational graph auto-encoder
We adapt the variational graph auto-encoder from Kipf to the bipartite case by considering two graph convolutional networks (GCN), one for each node type. We consider a biadjacency matrix $B$ of size $n_{1}\times n_{2}$ representing our bipartite graph. For all $i$ and $j$, $B_{i,j}\in\{0,1\}$ denotes the absence or the presence of interaction between the $i^{th}$ node of the first group and the $j^{th}$ node of the second group. Let
@@ 2.1.2 Decoder
Following rubin-delanchyStatisticalInterpretationSpectral2021, we decide to use as a decoder the generalised random dot product
@@
$p(B_{i,j}|Z_{1i},Z_{2j})=sigmoid(Z_{1i}^{\top}\mathbf{I}_{D_{+},D_{-}}Z_{2j})$ where $sigmoid:x\mapsto\frac{1}{1+e^{-x}}$ and $\mathbf{I}_{D_{+},D_{-}}$ is a diagonal matrix with $D_{+}$ ones followed by $D_{-}$ minus ones on its diagonal, such as $D_{+}+D_{-}=D$. We selected this decoder because its versatility covers a large family of graph structures. The full auto-encoder can be summarized as
@@ 2.2.1 Definition
Let $X$ and $Y$ two random variables in $\mathcal{X}$ and $\mathcal{Y}$ and $(X,Y)$ be the joint probability distribution. Let $\mathcal{F}$ and $\mathcal{G}$ be the RKHS on $\mathcal{X}$ and $\mathcal{Y}$ with their associated kernel $K:\mathcal{X}\times\mathcal{X}\to\mathbb{R}$ and $L:\mathcal{Y}\times\mathcal{Y}\to\mathbb{R}$. grettonMeasuringStatisticalDependence2005a define the HSIC as the norm of the cross-variance operator between the distribution in the RKHS:
[…]
@@
[…]
@@ 2.2.2 Estimation
Under the assumption that $X$ and $Y$ are independent, it has been proved that the distribution of $n\times\widehat{HSIC}$ can be asymptotically approximated by a Gamma distribution (gretton_gamma):
@@
During the learning phase, using the HSIC as a penalty between the coordinates in the latent space and the protected variable will assure that the latent space is as much as possible independent of the protected variable. At the end of the learning, we can check the independence by calculating the p-value of the test and compare it to the desired significance level. An example in the linear case is available in A.2.
@@ 2.3 Estimation in high dimension
Calculating $\widehat{HSIC}$ requires to compute the $n\times n$ Gram matrix and can be time-consuming for large value of $n$. As a substitute, we can use Random Fourier Features (RFF) (rahimiRandomFeaturesLargeScale2007). RFF for learning has been used in (louizosVariationalFairAutoencoder2017) to minimize the maximum mean discrepancy (MMD) between two distributions, in a setting where the protected variable have two possible values. Using the Gaussian kernel, minimizing the MMD is equivalent to matching all the moments of the two distribution, making the outcome of the encoder fair. In our settings, the HSIC can be seen as the MMD between the joint distribution $(X,Y)$ and the product of the distribution $X$ and $Y$. If the joint distribution is close to the product of the two distribution then the $X$ and $Y$ will behave as they are independent.
[…]
$\psi_{X}(x_{i})=\sqrt{\frac{2}{h}}cos(\omega x_{i}^{\top}+b)\in\mathbb{R}^{h}.$
The main idea behind this RFF representation is that
$K(x_{i},x_{j})\approx\langle\psi_{X}(x_{i}),\psi_{X}(x_{j})\rangle.$
This property will allow the computation to be much faster (see B.4), while having a small error term (sutherlandErrorRandomFourier2015). If we also define $\psi_{Y}$ the as RFF of $Y$, then finally we can estimate the RFF HSIC (zhangLargescaleKernelMethods2018) as done in quadriantoDiscoveringFairRepresentations2019:
[…]
The RFF HSIC can then be computed accurately with complexity $O(h^{2})$. Several other methods of estimation for HSIC exists for large-scale problem (zhangLargescaleKernelMethods2018). For the rest of this paper, $h=100$ will provide enough accuracy.
2.4 Bipartite and fair graph variational auto-encoder
Suppose our goal is to construct the latent representation $Z_{1}$ of the bipartite variational auto-encoder such as it is independent of a protected variable denoted by $S$. We optimize our parameters $(W_{v,i})_{1\leq v,i\leq 2}$ as defined in Section 2.1.1 to minimize a compromise between the variational lower bound of the auto-encoder and the HSIC between $\mu_{1}$ and the protected variable $S$. Even if the reconstruction would be penalized, this would yield a latent-space $\tilde{Z_{1}}\sim\mathcal{N}(\mu_{1},diag(\sigma^{2}_{1}))$ independent of $S$. The complete loss of this auto-encoder can be written as: